# Dyonic Taub-NUT-AdS Black Branes: Thermodynamics and Phase Diagrams

### v1

@@ Abstract
Motivated by the recent developments in the thermodynamics of Taub-NUT spaces and the absence of Misner strings in Taub-NUT solutions with flat horizons, we investigated the phase structure of dyonic Taub-NUT solutions. We follow the treatment proposed in arXiv:2206.09124 and arXiv:2304.06705 to introduce the nut parameter as a conserved charge to the first law. Although the calculated quantities satisfy the first law, we have found a larger class of charges that satisfy the first law and depend on some arbitrary parameter which we call $\alpha$. We choose to describe phase diagrams as NUT parameter-Temperature graphs to show borders of big and small black hole phases. We study the phase structure of these spaces in a mixed ensemble (i.e., we fix the electric potential, the nut parameter, and the magnetic charge), which we classify into different cases depending on the value of $\alpha$. In some of these cases we have first-order phase transitions that end with critical points. These classes could include up to four critical points, again, depending on $\alpha$ and the other quantities.
@@ 1 Introduction
One of the most attractive features of anti-de Sitter space solutions is the possibility of having positive specific heat [1]. This reflects the thermal stability of these AdS solutions in contrast to flat space solutions with negative specific heat. In their work, Hawking and Page [1] were able to show the existence of a phase transition between the black hole and the pure thermal AdS phases. Later, this was interpreted as the transition between the confinement and deconfinement phases of the dual gauge theory in the context of AdS/CFT correspondence [2]. After a year, this was extended to the AdS charged black hole [3, 4] where the authors found first-order phase transitions ending at a critical point, a very similar behavior to the well known liquid-gas phase transition.
@@
Indeed, Working in AdS spaces enlarges the spectrum of various types of black holes, since they come with three horizon topologies; spherical, flat, or hyperbolic. Therefore, it is interesting to study the thermodynamics of AdS black holes with a flat horizon, which was investigated in [5, 6, 7]. Furthermore, their extended thermodynamics was studied in [8] as well. However, these investigations showed no sign of first-order phase transitions or critical behavior since the black hole horizon temperature varies monotonically as a function of the horizon radius! This is why some authors tried to use unconventional methods to obtain a system with a non-trivial phase diagram as in [9, 10, 11, 12]. In [9, 10, 11], for example, the authors introduced a complex scalar field that gave rise to a superconductor phase diagram with a second-order phase transition between a black brane with scalar hair at low temperature and a black brane with no hair at high temperature. In [12], on the other hand, the authors achieved phase transition and critical behavior by introducing higher curvature terms with certain topological features to the gravitational action.
@@
One of the attractive classes of solutions studied in this duality is the AdS-Taub-Bolt/NUT black holes in four and higher dimensions [13, 14, 15, 16]. This class of solutions is a generalization to the Schwarzschild-AdS space with one additional parameter, the nut charge ”n”. An important feature of these solutions is that they possess string-like singularities known as Misner strings[17]. These strings are analogous to the Dirac string in electromagnetism and can be removed by imposing certain periodicity conditions, namely; the inverse temperature set to be $\beta=8\pi n$. In [13, 18, 14, 19, 20] the thermodynamics was studied and the authors revealed unusual features of these solutions. For example, the entropy is not the area of the horizon, even not always positive, and the horizon temperature and radius are fixed by the nut parameter n. Another unusual property of these solutions is that there are no additional terms in the first law due to the above relation between the nut charge and the horizon radius.
Motivated by some recent suggestions on the Taub-NUT thermodynamics [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] several authors tried to probe the possibility of relaxing the above periodicity condition, leaving the Misner string visible and allowing the NUT charge to appear in the first law as an independent charge with its conjugate potential.
In this work we study the regular thermodynamics of dyonic black branes with a non-vanishing NUT charge using the treatment introduced in [30, 33] where the authors introduced the NUT charge as a conserved charge which contributes to the first law with an extra work-term. Also, in [33], the authors found mass, angular momentum, NUT, electric, and magnetic charge densities distributed along Misner string when studying the dyonic Taub-NUT black hole solution. In fact, relaxing the periodicity condition was first considered by Bonner in [36] where he showed that the string singularity is a source of angular momentum.
@@
In this study, we closely follow the approach outlined in [34], but with two major differences. Firstly, instead of focusing on black holes with spherical horizons, we will be examining black holes with flat horizons, also known as ”black branes”. Secondly, while authors in [34] studied the extended thermodynamics of black holes, which includes considerations such as varying the cosmological constant as a thermodynamic variable, our study will focus on the regular thermodynamics of black branes. We show that allowing the dyonic black brane to have a non-vanishing NUT charge while relaxing the Misner periodicity condition results in a consistent thermodynamics where the Gibbs-Duhem relation and the first law are both satisfied, the entropy is related to the horizon area, and the temperature goes to that of the NUT-less case as $n\rightarrow 0$. We also show that the phase diagram becomes nontrivial in this case, with first-order phase transitions, continuous phase transitions, and critical behavior taking place.
@@
The paper is organized as follows; in section (2) we revisit the NUT-less dyonic black brane showing that its phase diagram is indeed trivial with no signs of phase transitions or critical behavior. In section (3) we study the Taub-NUT dyonic black brane, studying the regular thermodynamics of the mixed ensemble of the solution by calculating its thermodynamic quantities and showing that they alongside an arbitrary parameter $\alpha$ satisfy the first law and the Gibbs-Duhem relation in subsection (3.1), In (3.2) we divide the general case into sub-cases depending on the values of the thermodynamic quantities and studying the phase transitions and the phase diagram for each case. Finally in section (4) we discuss our treatment and present our conclusion while outlining possible extensions and future directions.
@@ 2 NUT-less Dyonic AdS Black Brane
Where the one form $B$ is the solution for $dB\,=\,*F$, where $*F$ is the Hodge dual of $F$, given by
@@ 2.1 Thermodynamics
[…]
@@
$\left(\frac{\partial{\Omega}}{\partial{T}}\right)_{\phi_{e},Q_{m}}=-S,\quad\left(\frac{\partial{\Omega}}{\partial{\phi_{e}}}\right)_{T,Q_{m}}=-Q_{e},\quad\left(\frac{\partial{\Omega}}{\partial{Q_{m}}}\right)_{T,\phi_{e}}=\phi_{m}.$ (23)
@@
$\left(\frac{\partial{\mathfrak{M}}}{\partial{S}}\right)_{Q_{e},Q_{m}}=T,\quad\left(\frac{\partial{\mathfrak{M}}}{\partial{Q_{e}}}\right)_{S,Q_{m}}=\phi_{e},\quad\left(\frac{\partial{\mathfrak{M}}}{\partial{Q_{m}}}\right)_{S,Q_{e}}=\phi_{m}.$ (26)
@@ 3 Taub-NUT Dyonic AdS Black Brane
$Q_{e}=\frac{1}{4\pi\sigma}\int_{\partial\Sigma}*F,\quad\,Q_{m}=-\frac{1}{4\pi\sigma}\int_{\partial\Sigma}F$ (35)
@@
$=p+2n\phi_{e}.$ (37)
In fact, if we choose the boundary at any generic radial distance with $r\,>\,r_{h}$, we will get the same results for the electric and magnetic charge densities. This is a very interesting result. Unlike the NUT-less case where there was no flux through the constant $x$ hypersurface, we get here an $r$-dependent electric and magnetic charge densities if we considered the surfaces of constant $r$ for any generic radial distance with $r>r_{h}$. This is a consequence of the flux leaking through the constant $x$ surface.
We can also calculate the conserved NUT charge density, $N$, as the dual to the mass density by the generalized Komar integral [41, 21]
@@ 3.1 Thermodynamics
[…]
@@
$\left(\frac{\partial{\Omega}}{\partial{T}}\right)_{\phi_{e},\widetilde{Q}_{m},n}=-S,\quad\left(\frac{\partial{\Omega}}{\partial{\phi_{e}}}\right)_{T,\widetilde{Q}_{m},n}=-\widetilde{Q}_{e},$
$\left(\frac{\partial{\Omega}}{\partial{\widetilde{Q}_{m}}}\right)_{T,\phi_{e},n}=\phi_{m},\quad\left(\frac{\partial{\Omega}}{\partial{n}}\right)_{T,\phi_{e},\widetilde{Q}_{m}}=\phi_{n},$ (45)
@@
$\left(\frac{\partial{\mathfrak{M}}}{\partial{S}}\right)_{\widetilde{Q}_{e},\widetilde{Q}_{m},n}=T,\quad\left(\frac{\partial{\mathfrak{M}}}{\partial{\widetilde{Q}_{e}}}\right)_{S,\widetilde{Q}_{m},n}=\phi_{e},$
$\left(\frac{\partial{\mathfrak{M}}}{\partial{\widetilde{Q}_{m}}}\right)_{S,\widetilde{Q}_{e},n}=\phi_{m},\quad\left(\frac{\partial{\mathfrak{M}}}{\partial{n}}\right)_{S,\widetilde{Q}_{e},\widetilde{Q}_{m}}=\phi_{n}.$ (50)
@@
In the next subsection, we will show that the phase diagram of the NUTty solution contains 0, 1, 2, or 4 critical points, in contrast to Nutless case. The number of critical points depends on the values of $\alpha$, $\widetilde{Q}_{m}$, and $n$. The other free parameters of the solution, $l$ and $\phi_{e}$, only scale the phase diagram by changing the values of $\alpha$, $\widetilde{Q}_{m}$, and $n$ where critical behavior occurs. Therefore, we can choose the values of $l$ and $\phi_{e}$ arbitrarily without losing any generality.
3.2 Phase Transitions and Critical Points
Since the Dyonic Nutless case has no phase transitions to compare it to the present cases, we find it convenient to construct the phase diagram as a $n-T$ diagram. The values of $\alpha$ and $\widetilde{Q}_{m}$ will affect and characterize these phase diagrams.
Before starting our analysis, we can notice that the signs of $\alpha$ and $\widetilde{Q}_{m}$ only affect the term $\left(\widetilde{Q}_{m}-\alpha n\phi_{e}\right)^{2}$. If both $\alpha$ and $\widetilde{Q}_{m}$ switch signs, the system is unaffected, but if only one of $\alpha$ or $\widetilde{Q}_{m}$ switches sign, we will get a system where $n\longrightarrow-n$. In the following analysis we will use positive values for $\alpha$ and $\widetilde{Q}_{m}$ knowing that we are not losing any generality.

### v2

@@ Abstract
Motivated by the recent developments in the thermodynamics of Taub-NUT spaces and the absence of Misner strings in Taub-NUT solutions with flat horizons, we study the phase structure of the dyonic Taub-NUT solutions with Lorentzian signature. We follow the treatment proposed in arXiv:2206.09124 and arXiv:2304.06705 to introduce the NUT parameter as a conserved charge to the first law. Although the calculated quantities satisfy the first law, we have found a larger class of charges that satisfy the first law and depend on some arbitrary parameter $\alpha$. We choose to describe phase diagrams as NUT parameter-Temperature graphs to show borders of big and small black brane phases. We study the phase structure of these spaces in a mixed ensemble (i.e., fixed electric potential, NUT charge, and magnetic charge), which we classify into different cases depending on the value of $\alpha$ and the other quantities. Most of these cases not only have first-order phase transitions, but also, end at critical points. Some of these cases include up to four critical points, depending on the value of $\alpha$ and the other quantities.
@@ 1 Introduction
One of the most attractive features of anti-de Sitter space solutions is their positive specific heat [35]. This feature reflects thermal stability of the system in contrast with the known negative specific heat of asymptotically flat spacetimes. In their seminal analysis of Schwarzschild-AdS black holes, Hawking and Page [35] were able to show the possibility of phase transitions between the black hole and the pure thermal AdS phases. Upon discovering the AdS/CFT correspondence [43, 53], this transition was interpreted as the confinement and deconfinement phase transition of the dual gauge field theory.
In short, the AdS/CFT correspondence is a duality that relates an (n+1n+1)–dimensional theory of gravity on anti–de Sitter (AdS) spacetime to a gauge/conformal field theory (CFT) in the boundary that has $n$ dimensions. In this duality, there is a dictionary linking every quantity in the bulk (AdS-gravity) to another in the boundary (quantum field theory). For example, a mass in Schwarzschild-AdS is linked to a finite temperature on the boundary field theory.
In 1999 Hawking-Page analysis was extended to the AdS charged black hole in [16, 17], where the authors found first-order phase transitions, which end at a critical point. This transition is analogous to the well-known liquid-gas phase transition, or Van der Waals fluid. Later, many authors studied the phase structure of various charged AdS black holes, either by considering a fixed cosmological constant (Check for example, Refs. [23, 24, 34, 33]), or allowing it to vary, which is known as extended phase space thermodynamics (Check for example, Refs. [41, 27, 38, 51]).
@@
Working in AdS spaces enlarges the spectrum of various types of black holes, since they come with three horizon topologies; spherical, flat, or hyperbolic. Therefore, it is interesting to study the thermodynamics of AdS black holes with a flat horizon, which was investigated in [52, 9, 13]. Furthermore, their extended thermodynamics was studied in [27] as well. However, these investigations showed no sign of first-order phase transitions or critical behavior since the black hole horizon temperature varies monotonically as a function of the horizon radius! However, several authors were able to show the existence of phase transition and critical behavior for these solutions by adding more exotic ingredients to the setup [34, 33, 50, 14, 27, 49, 38, 39]. In [33, 34, 39], for example, the authors introduced a complex scalar field that gave rise to a superconductor phase diagram with a second-order phase transition between a black brane with a scalar hair at low temperature and a black brane with no hair at high temperature. Alternatively, in [38], the authors achieved phase transition and critical behavior by introducing higher curvature terms with certain topological features to the gravitational action.
@@
One of the attractive AdS solutions studied in the above contexts is the AdS-Taub-Bolt/NUT black holes in four and higher dimensions [18, 28, 48, 3]. This class of solutions is a generalization to the Schwarzschild-AdS space with one additional parameter, the NUT parameter ”n”. An important feature of these solutions is that they possess string-like singularities known as Misner strings [47]. These strings are analogous to the Dirac string in electromagnetism and can be removed by imposing certain periodicity conditions, namely; the inverse temperature set to be $\beta=8\pi n$.
The Taub-NUT solution is often considered to be a gravitaional dyon [25], where the dual to the Komar mass is non-vanishing. This dual quantity defines the NUT charge $N$, interpreted as the gravitational analogue of magnetic charge. For Taub-NUT-AdS solution, the conserved charge $N$ can be calculated by taking the dual of the generalized Komar integral [40, 11]
$N=\frac{1}{4\pi}\int_{\partial\Sigma}\mathrm{d}\xi-2\Lambda(*\omega)=n\left(k-\frac{4n^{2}}{l^{2}}\right),$
where $k$ is the curvature index , $k=0$ for the flat-horizon, and $k=\pm 1$ for the spherical and hyperbolic horizons, respectively. Note that for regular thermodynamics where the AdS radius, $l$, is fixed, it follows that the NUT parameter $n$ is also a conserved quantity.
But what does the NUT parameter on the boundary correspond to? In the early works [18, 28, 45], where the authors used Euclidean signature, the NUT parameter ”n” was interpreted as the squashing parameter controlling the deformation of the boundary manifold, i.e., the 3-sphere. A deeper interpretation appeared after studying the Taub-NUT solutions with Lorentzian signatures in the context of fluid/gravity correspondence[42, 15]. It provided a link between the bulk properties and the hydrodynamics of its boundary quantum field theory. The NUT parameter was interpreted as a measure of vorticity in the flux of the dual conformal fluid [42]. It showed that a Taub-NUT-AdS solution is dual to a fluid characterized by a constant vortex flow [42, 15] with a Dirac delta source, generated by the bulk Misner string.
Many authors studied Taub-NUT solutions with Euclidean signature while enforcing the Misner periodicity condition [18, 45, 28, 20, 2]. They found that the thermodynamics of these solutions revealed unusual features. For example, the entropy is not the area of the horizon, even not always positive, and the horizon temperature and radius are fixed by the nut parameter n. Furthermore, in the $n\rightarrow 0$ limit, one expects to retrieve the known Schwarzschild-AdS temperature, instead, one gets a diverging temperature! Another unusual property of these solutions arises from the observation that the NUT charge, being a conserved quantity, should contribute an additional work term in the first law. However, no such contribution is present due to the relation between the nut charge and the horizon radius.
Motivated by the above issues and some recent developments several authors [11, 12, 19, 37, 7, 54, 8, 26, 29, 4, 44, 1, 5, 6, 51] tried to probe the possibility of relaxing the periodicity condition, leaving the Misner string visible and allowing the NUT charge to appear in the first law as an independent charge (Please check the ”Introduction” section of [5] for a quick review). This new approach was first adopted by Bonner in [10] where he showed that the string singularity is a source of angular momentum. Also in [5], the authors found mass, angular momentum, NUT, electric, and magnetic charge densities distributed along Misner string when studying the dyonic Taub-NUT black hole solution. Moreover, other authors showed in [46, 21, 22] that the spacetime can be maximally extended and geodesically complete by relaxing the Misner periodicity condition without any causal pathologies.
@@
In this work we study the thermodynamics, with fixed cosmological constant, of the dyonic black branes with a non-vanishing NUT charge. For Taub–NUT solutions with flat horizons, the absence of a Misner string removes the need to impose the Misner periodicity condition [18]. Consequently, the NUT charge remains as an independent conserved quantity, contributing to the first law with an additional work term. Our work will closely follow the approach outlined in [5, 6], but with two major differences. Firstly, instead of focusing on black holes with spherical horizons, we will be examining black holes with flat horizons, also known as ”black branes”. Secondly, while the authors in [6] studied the extended thermodynamics of black holes, i.e., considers varying the cosmological constant as a thermodynamic variable, our study will focus on regular thermodynamics of black branes. We show that allowing the dyonic black brane to have a non-vanishing NUT charge while relaxing the Misner periodicity condition results in a consistent thermodynamics where the Gibbs-Duhem relation and the first law are both satisfied, the entropy is related to the horizon area, and the temperature goes to that of the NUT-less case as $n\rightarrow 0$. We also show that the phase diagram becomes nontrivial in this case, with first-order phase transitions ending at critical points.
@@
The paper is organized as follows; in section (2) we revisit the NUT-less dyonic black brane showing that its phase diagram is indeed trivial with no signs of phase transitions or critical behavior. In section (3) we study the Taub-NUT dyonic black brane, studying the regular thermodynamics of the mixed ensemble of the solution by calculating its thermodynamic quantities and showing that they satisfy the first law and the Gibbs-Duhem relation in subsection (3.1), In (3.2) we divide the general case into sub-cases depending on the values of the thermodynamic quantities and studying the phase transitions and the phase diagram for each case. Finally in section (4) we discuss our treatment and present our conclusion while outlining possible extensions and future directions.
@@ 2 NUT-less Dyonic AdS Black Brane
Where the one form $B$ is the solution for dB=FdB\,=\,*F, where F*F is the Hodge dual of $F$, given by
@@ 2.1 Thermodynamics
$T_{h}\,=\,\frac{f^{\prime}(r)}{4\,\pi}\biggr|_{r=r_{h}}\,=\,\frac{1}{\,4\,\pi\,r_{h}}\left(\frac{3\,r_{h}^{2}}{l^{2}}-\frac{Q_{m}^{2}}{r_{h}^{2}}-\phi_{e}^{2}\right).$ (14)
@@
$\left(\frac{\partial{\Omega}}{\partial{T}}\right)_{\phi_{e},Q_{m}}=-S,\qquad\left(\frac{\partial{\Omega}}{\partial{\phi_{e}}}\right)_{T,Q_{m}}=-Q_{e},\qquad\left(\frac{\partial{\Omega}}{\partial{Q_{m}}}\right)_{T,\phi_{e}}=\phi_{m}.$ (23)
@@
$\left(\frac{\partial{\mathfrak{M}}}{\partial{S}}\right)_{Q_{e},Q_{m}}=T,\qquad\left(\frac{\partial{\mathfrak{M}}}{\partial{Q_{e}}}\right)_{S,Q_{m}}=\phi_{e},\qquad\left(\frac{\partial{\mathfrak{M}}}{\partial{Q_{m}}}\right)_{S,Q_{e}}=\phi_{m}.$ (26)
@@ 3 Taub-NUT Dyonic AdS Black Brane
$Q_{e}=\frac{1}{4\pi\sigma}\int_{\partial\Sigma}*F,\qquad\,Q_{m}=-\frac{1}{4\pi\sigma}\int_{\partial\Sigma}F$ (35)
@@
$p+2n\phi_{e}.$ (37)
In fact, if we choose the boundary at any generic radial distance with $r\,>\,r_{h}$, we will get the same results for the electric and magnetic charge densities. This is a very interesting result. Unlike the NUT-less case where there was no flux through the constant $x$ hypersurface, we get here an $r$-dependent electric and magnetic charge densities if we only considered the surfaces of constant $r$ for any generic radial distance with $r>r_{h}$. This is a consequence of the flux leaking through the constant $x$ surface.
We can also calculate the conserved NUT charge density, $N$, as the dual to the mass density given by the generalized Komar integral [40, 11]
@@ 3.1 Thermodynamics
It is important to see that the charge densities calculated earlier, $Q_{e}=\left(q-2n\phi_{m}\right)$ and $Q_{m}\,=\,\left(p\,+\,2n\phi_{e}\right)$, do not satisfy the first law and the Gibbs-Duhem relation. In fact, if we set the electric charge density to be $Q_{e}$, the magnetic charge density that satisfies the first law and the Gibbs-Duhem relation is $p$. And if we set the magnetic charge density to $Q_{m}$, the electric charge that satisfies the two relations is nothing but $q$. In short, to satisfy these relations one of the charges must be the charge at infinity and the other one the charge at the horizon! This fact was previously reported by several authors [12, 19, 8, 4, 5, 6, 51]. Two of these authors [19, 5] showed that there exist an infinite family of electric and magnetic charge densities that satisfy both relations. The two charge densities can be parameterized by a real parameter $\alpha$, in the following forms:
@@
$\left(\frac{\partial{\Omega}}{\partial{T}}\right)_{\phi_{e},\widetilde{Q}_{m},n}=-S,\qquad\left(\frac{\partial{\Omega}}{\partial{\phi_{e}}}\right)_{T,\widetilde{Q}_{m},n}=-\widetilde{Q}_{e},$
$\left(\frac{\partial{\Omega}}{\partial{\widetilde{Q}_{m}}}\right)_{T,\phi_{e},n}=\phi_{m},\qquad\left(\frac{\partial{\Omega}}{\partial{n}}\right)_{T,\phi_{e},\widetilde{Q}_{m}}=\phi_{n},$ (45)
@@
$\left(\frac{\partial{\mathfrak{M}}}{\partial{S}}\right)_{\widetilde{Q}_{e},\widetilde{Q}_{m},n}=T,\qquad\left(\frac{\partial{\mathfrak{M}}}{\partial{\widetilde{Q}_{e}}}\right)_{S,\widetilde{Q}_{m},n}=\phi_{e},$
$\left(\frac{\partial{\mathfrak{M}}}{\partial{\widetilde{Q}_{m}}}\right)_{S,\widetilde{Q}_{e},n}=\phi_{m},\qquad\left(\frac{\partial{\mathfrak{M}}}{\partial{n}}\right)_{S,\widetilde{Q}_{e},\widetilde{Q}_{m}}=\phi_{n}.$ (50)
@@ 3.2 Phase Transitions and Critical Points
Before starting our analysis, we can notice that the signs of $\alpha$, $\widetilde{Q}_{m}$, and $\phi_{e}$ only affect the term $\left(\widetilde{Q}_{m}-\alpha n\phi_{e}\right)^{2}$. If two of the variables switch signs, the system is unaffected, but if all of them or only one them switches sign, we will get a system where $n\longrightarrow-n$. In the following analysis we will use positive values for $\alpha$, $\widetilde{Q}_{m}$, and $\phi_{e}$ knowing that we are not losing any generality.
Since the Dyonic NUTless case has no phase transitions to compare it to the present cases, we find it convenient to construct the phase diagram as a nTn-T diagram to highlight the role of the NUT charge in developing first order phase transitions and critical behaviors.
By allowing non-zero values for $n$, we aim to check if we will get a non-monotonic behavior for $T_{h}(r_{h})$, which indicates first order phases, and, if so, we require to get the critical points, which are the values of $n$ at which $T_{h}(r_{h})$ goes from being monotonic to non-monotonic. Doing so, we got 4 critical points in general
$n_{c_{1,2}}=-\alpha l\phi_{e}\mp\frac{1}{3}{\sqrt{3l\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}},$ (52)
corresponding to
$r_{h(c_{1,2})}=\frac{1}{3}{\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}+l\widetilde{Q}_{m}\right)\pm 3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}}\,$ (53)
and
$n_{c_{3,4}}=\alpha l\phi_{e}\mp\frac{1}{3}{\sqrt{3l\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}},$ (54)
corresponding to
$r_{h(c_{3,4})}=\frac{1}{3}{\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}-l\widetilde{Q}_{m}\right)\mp 3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}},$ (55)
Some of these expressions may become imaginary, which correspond to non-physical critical points. However, for each critical point, the same square-root term that determine the reality of $n_{c}$ also appear in $r_{h(c)}$. Thus, ensuring the reality of $r_{h(c)}$ automatically guarantees that $n_{c}$ is also real.
Assuming positive $\widetilde{Q}_{m}$ and $\phi_{e}$, one can see that while the first critical point is always physical, the other critical points come with restrictions on the thermodynamic quantities in order to physically exist.
  • Second critical point:
    For this critical point to exist, relation for $r_{h(c_{2})}$ imposes the following condition
    $\widetilde{Q}_{m}>\frac{\alpha l\phi^{2}}{\sqrt{3}}$ (56)
  • Third critical point:
    The conditions required for $r_{h}(c_{3})$ to be real are
    $\alpha>\frac{\sqrt{3}}{3}$ (57)
    and
    $\frac{\alpha l\phi^{2}}{\sqrt{3}}<\widetilde{Q}_{m}\leq\frac{1}{6}\left[l\phi^{2}\left(3\alpha^{2}+1\right)\right].$ (58)
  • Fourth critical point:
    Finally for the fourth critical point, $r_{h(c_{4})}$ is real when
    $\alpha\leq\frac{\sqrt{3}}{3},\qquad\text{and}\qquad\widetilde{Q}_{m}<\frac{\alpha l\phi^{2}}{\sqrt{3}},$ (59)
    or
    $\alpha>\frac{\sqrt{3}}{3},\qquad\text{and}\qquad\widetilde{Q}_{m}\leq\frac{1}{6}\left[l\phi^{2}\left(3\alpha^{2}+1\right)\right].$ (60)
From the above conditions we can see that the third critical point is always non-physical for $\alpha\leq\frac{\sqrt{3}}{3}$ regardless of the values of $\widetilde{Q}_{m}$ and $\phi_{e}$. Also, for $\alpha\leq\frac{\sqrt{3}}{3}$, the second and fourth critical points cannot be real simultaneously, implying that at most only two critical points can exist simultaneously in this regime. On the other hand, for $\alpha>\frac{\sqrt{3}}{3}$, all four critical points can be real simultaneously depending on the values of $\widetilde{Q}_{m}$ and $\phi_{e}$. Accordingly, we will split our analysis depending on the value of $\alpha$. Dependencies on the other thermodynamic quantities will be considered as sub-cases. In our analysis we chose to fix the value of $\phi_{e}$ and parametrize the sub-cases using $\widetilde{Q}_{m}$. The reverse, fixing the value of $\widetilde{Q}_{m}$ and parametrize the sub-cases with $\phi_{e}$, is also valid and will lead to the same results. Therefore, this choice will not lead to any loss of generality.